There is an equilateral triangle where each side is 2x-7, x+y-9 and y+5
Calculate the value of x and find the length of the sides of the triangle

Answer by ankor@dixie-net.com(15747)

There is an equilateral triangle where each side is 2x-7, x+y-9 and y+5
:
All side equal, therefore
2x - 7 = x + y - 9
2x - x - y = -9 + 7
x - y = -2
and
2x - 7 = y + 5
2x - y = 5 + 7
2x - y = 12
Subtract the 1st equation from the above equation
2x - y = 12
x - y = -2
---------------subtraction eliminates y
x = 14
find y
x - y = -2
14 - y = -2
-y = -2 -14
-y = -16
y = 16
:
Find the 3 sides, a, b, c
a = 2x-7
a = 2(14) - 7
a = 28 - 7
a = 21
:
b = x + y - 9
b = 14 + 16 - 9
b = 21
:
c = y + 5
c = 16 + 5
c = 21